Circular No. 5673

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

AN URSAE MAJORIS J. M. Bonnet-Bidaud, Centre d’Etudes Nucleaires de Saclay; M.

Mouchet, Observatoire de Meudon; and T. A. Somova and N. N. Somov, Special Astronomical Observatory, communicate: “Optical spectroscopic data obtained at the Russian 6-m telescope during 1991 Mar.-1992 Jan. show a significant phase shift (0.20 cycle) in the velocity curves of AN UMa with respect to the ephemeris period of Liebert et al. (1982, Ap.J. 254, 232), evidently due to an inaccuracy in the orbital/rotation period. From a reanalysis of all published spectroscopic data for the narrow-line components, we have determined a new ephemeris for the ‘blue-to-red gamma crossing time’: 1977 Feb. 16.5255 (± 0.0024 day) UT + 0.079 752 82 (± 0.000 000 04) E. With this new period, we find that all published photometric light curves are also correctly phased over more than 15 yr. The updated ephemeris eliminates, in particular, the phase discrepancy reported in x-rays; the 1984-85 EXOSAT light curves now show x-ray minimum near phase 0.0 (with the above ephemeris), compatible with previous SAS-3 data. This is evidence for the geometrical stability of the system. To the accuracy of the measurements, we find no evidence of a higher term of period variation (Ṗ < 10^{−10}) nor of any asynchronism in the system. Polarimetric observations are strongly encouraged to improve the accuracy of the revised period. We predict a linear polarimetric maximum at a present offset of 0.35 cycles with the previous (Liebert et al.) ephemeris.”

PERIODIC COMET SWIFT-TUTTLE (1992t) U. Fink and M. Hicks, University of Arizona, report: “We regret to be

forced to issue a retraction to our report of the split nuclei of P/Swift-Tuttle (cf. IAUC 5672). Though the secondary nuclei appeared quite strongly in our H_{2}O^{+} filter images, they have not shown up in any of our other filter images. After carefully checking our standard- and comparison-star exposures through our H_{2}O^{+} filter, we find similar faint ‘satellite’ images as on the exposures of the comet. It would seem that the secondary nuclei are a reflection in the optics and therefore spurious.”

Total visual magnitude estimates: Nov. 28.71, 4.8 (H. Lüthen, Dahlenburg, Germany, 20×60 binoculars); 30.28, 5.3 (A. Mizser, Budapest, Hungary, 20×60 binoculars); Dec. 4.68, 5.3 (L. T. Heen, Trondheim, Norway, 7×50 binoculars); 5.73, 4.9 (J.-C. Merlin, Le Creusot, France, 8×30 binoculars); 9.93, 5.0 (D. W. E. Green, Cambridge, MA, 12×50 binoculars).

1992 December 10 Daniel W. E. Green